# Spatiotemporal Density Correction of Multivariate Global Climate Model Projections using Deep Learning

### v2

@@ 2.6 Conditional density estimation using SPQR
2.7 Model evaluation metrics for SPCDE
@@
In the study presented in Section 3, we assess the efficacy of SPCDE based on five functionals of the joint distribution of TMAX and PRCP: the Wasserstein distance between the marginal densities of calibrated and observational data, lag-1 autocorrelations, pairwise spatial correlations, and their 0.95 quantiles. These are evaluated for both the observed (nClimGrid) and calibrated predictions. We also evaluated the proportion of zeros for PRCP, and the cross correlation between TMAX and PRCP. The Wasserstein distance (Kantorovich, 1960; Vasershtein, 1969) quantifies the overall fit of the marginal distributions of each variable in terms of the dissimilarity between the densities of the observed and calibrated data, whereas the proportion of zeros and the 0.95 quantiles focus specifically on the tail behavior of the distributions. The three correlation metrics ensure that the calibrated data maintains the dependence structure of the nClimGrid data.
@@
In addition to assessing goodness of fit of our density corrected data, we also compare it with QM and CCA. As described in Section 1, QM is a bias correction method which matches the quantiles of the observational data time series and the GCM time series. The asynchronous QM bias correction method sorts GCM and nClimGrid data vectors in ascending order (independently), and then fits linear regression to the sorted data. This can take the form of simple linear regression (e.g., Dettinger et al., 2004) or piecewise linear regression (Stoner et al., 2013). The QM methodology is applied separately for each month, grid cell, and variable (TMAX and PRCP). The goal is to obtain a linear map between the quantiles of the GCM and nClimGrid datasets. We can then provide GCM data for unobserved grid points or time periods as input to the fitted model and obtain bias-corrected TMAX and PRCP as output. The resulting values must subsequently be rearranged back into the original order to retrieve the final product, a continuous chronological time series of the bias-corrected values. Since QM in our example is applied to each variable separately, it fails to account for inter-variable dependencies (Bhowmik and Sankarasubramanian, 2019). This can distort relationships critical for understanding processes such as hydrological cycles (Seo et al., 2019). We use CCA (Bhowmik et al., 2017) as our other point of comparison. CCA employs multivariate sorting of variables based on their joint probability of occurrence such that their cross correlation is maintained. Afterwards, canonical correlation is used to bias correct variables simultaneously. Neither of these two approaches, however, explicitly models spatial or temporal dependence present in the data.
@@ 3.1 Overview
We implemented SPCDE for daily observational and GCM data from 1951–2014 on two separate $5\times 5$ grids as shown in Figure 1. We separated the data into a training period (1951–2000) and a testing period (2001–2014), and fitted the spatiotemporal SPCDE model to the training period data. The Vecchia neighbor sets are selected using max-min ordering, which can result in significant improvements over other coordinate based orderings (Guinness, 2018). We model each month’s data separately to account for seasonality. Additionally, since daily PRCP contains a large proportion of zeroes and SPQR does not work with semi-continuous data, we do not model PRCP directly. Instead, we model $\log(0.0001+\mbox{PRCP})$, a common approach when modeling precipitation data using continuous distributions (Rajagopalan and Lall, 1998). We found this transformation sufficient for the simultaneous estimation of the proportion of zeros in the data, combined with a post-processing step (described in Section 3.2), as well as high quantiles.
@@
For each month and at every location conditioned on its neighboring set of locations, we fitted SPQR models for TMAX and the transformed PRCP. Each SPQR model has several hyperparameters, which were chosen based on a grid search based on different hyperparameter values (not presented). We choose each SPQR model to have 2 hidden layers with 30 and 20 nodes, and 20 output knots for both variables. The batch size and learning rate is set to 100 and 0.001 respectively, and the SPQR model is run for 300 iterations. In each model, $20\%$ of the training data is used as a validation set, with early stopping enforced if the validation set loss fails to decrease for more than 5 iterations.
@@
The SPQR models (for each month, location, and variable) can be fit in parallel, since the Vecchia approximation decomposes the joint density of Y into independent univariate conditional densities. Predictions for every month can also be carried out independently, and therefore the projection/calibration step occurs in parallel. Within each month, however, predictions need to be made in the same sequence as the ordering of the Vecchia neighbor sets. Predictions for TMAX and PRCP were obtained for the testing period of 2001–2014 at each location using Algorithm 2.
3.2 Evaluating bias-corrected output based on SPCDE
@@
Figure 2: Density curves of climate model (Mod), observed (Obs), and calibrated (Cal) TMAX and PRCP data in the Southwest (SW) and Southeast (SE) average over location and month from 2001–2014. PRCP results are presented in the log scale.
@@
We first present results from density correction at both sets of locations for the validation period of 2001–2014. Figure 2 plots density curves of density corrected TMAX and log-scale PRCP for the SE and SW regions. Each curve pools the entire validation period data across all 25 locations, and we see that the calibrated TMAX and PRCP closely match the observed data. Note that the density plots for PRCP are displayed in the transformed log scale since visual inspection of the density plots are most informative in the scale that the data was analyzed. However, all remaining plots and metrics for PRCP are presented in the raw data scale. Fitting separate monthly models for TMAX captures the pronounced bimodal distribution in the SW region with different modes for observed and model data. On the other hand, there are substantial differences between the observed and model densities of PRCP at both tails of their respective distributions. For both variables, SPQR is able to successfully calibrate the model data to match the density of the observational data.
@@
Figure 3: Monthly 0.95 quantiles of TMAX and PRCP data in the Southwest (SW) and Southeast (SE), aggregated over 2001–2014. Each point corresponds to a particular month and location.
@@
Figure 4: Monthly Proportion of zeros in PRCP data for the Southwest (SW) and Southeast (SE), aggregated over 2001–2014. Each point corresponds to a particular month and location.
Figure 5: Monthly lag-1 autocorrelations of TMAX and PRCP data in the Southwest (SW) and Southeast (SE), aggregated over 2001–2014. Each point corresponds to a particular month and location.
Figure 6: Monthly cross-correlations between TMAX and PRCP in the Southwest (SW) and Southeast (SE), aggregated over 2001–2014. Each point corresponds to a particular month and location.
@@
Figures 34 plot the 0.95 quantiles of TMAX and PRCP, and the proportion of zeros for PRCP, respectively. Each panel displays the effect of density correction by comparing it against the observational data for every location and month. For the upper 0.95 quantiles, the uncalibrated model values for TMAX have noticeable bias, something that can be visually verified in Figure 2. The calibrated quantiles virtually eliminates this bias while also having lower uncertainty than the uncalibrated data. The density correction also improves upper quantile estimation for PRCP. However, it is unable to fully adjust for the bias, and there is noticeably more uncertainty in the PRCP estimates than TMAX. Part of this bias is attributable to the proportion of zeros in PRCP data, which varies significantly by season. We therefore treat model precipitation below $0.001$ mm as zero precipitation, since it is not measurable by physical instruments like rain gauges. This was also done for the SPQR estimates since the continuous fitted distribution for PRCP will never produce exact zeros. The uncalibrated model data (red) in Figure 4 severely underestimates the proportion of zeros, due to a well known phenomenon called the drizzle effect (see, e.g., Levey and Sankarasubramanian, 2024), where numerical climate models for precipitation very rarely generate zero precipitation values, instead usually having very small but positive values. The calibrated density of PRCP is able to accurately represent the proportion of zeros that are observed in the real data, indicating that our log-transform for modeling PRCP is adequate for this application.
@@
Figure 7: Annual spatial correlations of TMAX and PRCP data in the Southwest (SW) and Southeast (SE), aggregated over 2001–2014. Each point corresponds to a pair of locations.
@@
Figures 56 plot the effect of density correction on the autocorrelations for each variable, and the cross correlation between them, respectively. As in the previous two plots, points correspond to month and location specific estimates. In Figure 5, SPCDE is unable to improve the autocorrelation for SW TMAX from its model values, but for the remaining three variables, the calibrated data has lower bias and variability in its estimates of autocorrelation compared to the uncalibrated model data. For the cross correlations in Figure 6, the uncalibrated data has positive bias in the cross correlations for both regions. Density correction was able to virtually eliminate the bias and provided highly accurate and precise estimates of the cross correlation. This is in part because the conditional density of PRCP depends on both current and past TMAX values, and therefore SPQR models the dependence between the two variables.
Finally, Figure 7 shows pairwise spatial correlations for the calibrated and uncalibrated data. Each panel contains values for 300 pairs of locations. Values are concentrated around the diagonal, suggesting the the calibrated data has similar spatial correlation compared to the observed data. We do notice more uncertainty in the lower end of TMAX correlation values in the uncalibrated data, that the density correction is not able to entirely address, but overall the calibrated data has lower bias and uncertainty compared to the uncalibrated model data.
3.3 Comparison of SPCDE with CCA and QM
Table 1: Out-of-sample prediction accuracy in the density correction of TMAX and PRCP based on a spatiotemporal SPCDE model, compared with canonical correlation analysis (CCA) and quantile mapping (QM). Other than the Wasserstein distance metric, the remaining values correspond to the RMSE between observed and calibrated data for 2001–2014. Metrics are averaged across spatial locations (pairs of locations for spatial correlation). Best (smallest) values for each combination of region, variable, and metric are highlighted.

### v3

@@ 2.5 Conditional density estimation using SPQR
As previously mentioned, SPQR is a method to approximate continuous density functions. However, daily PRCP contains a large proportion of zeroes making it a semi-continuous variable. Therefore, instead of modeling PRCP directly, we model $\log(0.0001+\mbox{PRCP})$, a common approach when modeling precipitation data using continuous distributions (Rajagopalan and Lall, 1998). We found this transformation adequate for estimating the proportion of zeros in the data, combined with a post-processing step (described in Section 4.2). However, SPQR can underestimate high quantiles of heavy-tailed distributions, since the M-splines are bounded and therefore not amenable to extrapolation. We noted this issue in our numerical experiments and in the discussion section.
2.6 Model evaluation metrics for SPECD
@@
In the studiers presented in Sections 34, we assess the efficacy of SPECD based on three functionals each of the marginal distributions and the dependence structure of calibrated PRCP and TMAX. We evaluate the Wasserstein distance (Kantorovich, 1960; Vasershtein, 1969) between the marginal densities of the calibrated and observational data, which quantifies the goodness-of-fit for each variable. Similarly, we also evaluate the proportion of zeros for PRCP and the 0.95 quantile for PRCP and TMAX. These two quantities focus on the tail behavior of the two variables. For assessing whether SPECD can preserve the dependence structure within nClimGrid, we evaluate pairwise spatial correlations for both variables, and the cross correlation between PRCP and TMAX. For all metrics other than the Wasserstein distance, we report the mean absolute error (MAE) between the observed and the calibrated data.
@@
In addition to assessing goodness of fit of our density corrected data, we also compare it with QM and CCA. As described in Section 1, QM is a bias correction method which matches the quantiles of the observational data time series and the GCM time series. The asynchronous QM bias correction method sorts GCM and nClimGrid data vectors in ascending order (independently), and then fits linear regression to the sorted data. This can take the form of simple linear regression (e.g., Dettinger et al., 2004) or piecewise linear regression (Stoner et al., 2013). The QM methodology is applied separately for each month, grid cell, and variable (TMAX and PRCP). The goal is to obtain a linear map between the quantiles of the GCM and nClimGrid datasets. We can then provide GCM data for unobserved grid points or time periods as input to the fitted model and obtain bias-corrected TMAX and PRCP as output. Since QM in our example is applied to each variable separately, it fails to account for inter-variable dependencies (Bhowmik and Sankarasubramanian, 2019). This can distort relationships critical for understanding processes such as hydrological cycles (Seo et al., 2019). We use CCA (Bhowmik et al., 2017) as our other point of comparison. CCA employs multivariate sorting of variables based on their joint probability of occurrence such that their cross correlation is maintained. Afterwards, canonical correlation is used to bias correct variables simultaneously. Both CCA and QM are applied independently to each grid cell, and therefore do not account for spatial correlation in the process of bias correction. Similarly, the asynchronous nature of both algorithms lead to temporal dependence is not specifically accounted for. Both methods have to contend with a well known phenomenon in GCMs called the drizzle effect (see, e.g., Levey and Sankarasubramanian, 2024), where numerical climate models for precipitation very rarely generate zero precipitation values, instead usually having very small but positive values. QM and CCA address this by applying a wet-day fraction correction in the form of a random offset added to the zero precipitation values in the observational data (Pierce et al., 2015; Wootten et al., 2021). The offset for each zero precipitation value is drawn independently from a Uniform(0.001,0.1)(0.001,0.1) distribution, leading to unique non-zero precipitation values. After bias corrections, precipitation values smaller than $0.01$ mm is set to zero.
3 Numerical Study
3.1 Overview
We study the efficacy of SPECD using a numerical study whose data generating process approximates our application. Since we will be fitting SPECD for each month separately, the numerical study focuses on modeling a single 30-day month’s TMAX and PRCP data, available for a period of 64 years. The first 50 years are treated as training data and is used to fit the model. Density correction is carried out on all 64 years of data, with the first 50 years used to benchmark in-sample performance while the last 14 years used to measure out-of-sample performance. Calibrating the data that the model is trained on is analogous to density correction of historical data. In contrast, out-of-sample performance gauges the method’s ability for calibrating future projections, provided the distribution of the GCM data does not change between the two time periods.
Figure 3: Density curves of climate model (Mod) and observed (Obs) TMAX and PRCP data, based on a single replicate from the simulation study. PRCP curve is presented on the log scale.
Data generating process: We generate the data for $64\times 30$ days at $m=25$ locations $\{1,2,3,4,5\}^{2}$, i.e., on a $5\times 5$ grid. For spatial location $i$ and replicate $t$, there are four responses, meant to represent (1) observed temperature, (2) observed precipitation, (3) model temperature and (4) model precipitation. The $4m$ outcomes for each replicate are generated from multivariate skew-t distribution. The location parameter is $(1/m,...,m/m)\otimes(2,3,1,2)$ so that each outcome has a different spatial trend. The scale matrix is $\Omega\otimes R$ where $\Omega$ is the spatial correlation matrix whose $(i,j)^{th}$ element is equal to $\exp(-|i-j|/2)$ and $R$ is the $4\times 4$ correlation matrix with elements $R_{12}=-0.8$, $R_{34}=-0.4$ and all other off-diagonal elements have a magnitude of 0.5. The strength of dependence between temperature and precipitation is stronger in the observations than the model under this setup. The degrees of freedom is set to 20; the skewness is 0 for both temperature responses, 100 for observed precipitation, and 10 for model precipitation. This results in non-Gaussian data with more skewness in the observations than the model. After generating the replications from the multivariate skewed-t distribution, we apply a spatial smoother (Gaussian kernel with bandwidth 2) to variables representing model output so that the spatial dependence is stronger in model output variables.
Finally, we rescale both TMAX and PRCP to ensure their scales are similar to what we observe in the application. The observational TMAX is scaled using a min-max transform to be between 255285K255-285\,K, while the model TMAX is scaled to be between 250280K250-280\,K. For observational PRCP, we shift the center of the distribution to the $75^{th}$ quantile, and threshold everything below that to be 0. A similar transformation is carried out for the model PRCP, but using the median value. As a consequence, $75\%$ of the observational PRCP is 0, and half of the model PRCP is 0, reflecting real world scenarios where GCMs have far fewer zeros in the data compared to the observational record. Figure 3 plots the density curve for the 64 years of data based on one simulated dataset, pooled over the 25 grid cells. We see that the generated process emulates a hotter and drier month in the observational record compared to the climate model output.
Models and testing criteria: We compare four different bias/density correction approaches based on the metrics laid out in Section 2.6. The first two are classical QM and CCA. The third and fourth are variants of SPECD, corresponding to Algorithms 1 and 2, respectively. SPECD1, described in Section 2.3, models all locations independent of each other, but considers the cross correlation between TMAX and PRCP. SPECD2 is described in Section 2.4 and considers both spatial correlation within each variable and cross correlation between the variables. The two variants of SPECD aim to disambiguate potential improvements in calibration due to SPQR from those due to the spatial information encoded via the Vecchia approximation.
3.2 Results
(a) Simulation study results for training period of 1951–2000.
(b) Simulation study results for validation period of 2001–2014.
Figure 4: Comparison of QM and CCE against two SPECD methods without (SPECD1) and with (SPECD2) spatial correlation. The left column reports Wasserstein distance; the remaining columns report MAE. Lower values indicate greater accuracy.
Figure 4 contains boxplots of the metrics detailed in Section 2.6 based on the 100 datasets. The top rows in Figures 4(a) and 4(b) each contain three metrics for TMAX - the Wasserstein distance between the true and calibrated densities (tmax-wasdist), and MAEs for the $95^{th}$ quantile (tmax-quantile) and the spatial correlation between the grid cells (tmax-spatcorr). The final plot reports the MAE of the cross correlations between the variables (crosscorr). The bottom rows each contain four panels of PRCP metrics - the Wasserstein distance between the true and calibrated densities (prcp-wasdist), and MAEs for the $95^{th}$ quantile (prcp-quantile), the spatial correlation (prcp-spatcorr), and the proportion of zeros (prcp-propzero). Each individual metric is evaluated based on the entire data - i.e., 50 years for Figure 4(a) and 14 years for Figure 4(b). Additionally, for all metrics except for the spatial correlation, the metrics are also pooled over the 25 locations.
We begin by noting that Figure 4(a) has smaller values overall than Figure 4(b), confirming that in-sample performance is better than out-of-sample performance. There is also increased variability in Figure 4(b), reflecting higher uncertainty in out-of-sample estimates. However, the two figures demonstrate broadly similar patterns otherwise. For the three TMAX metrics, both SPECD approaches outperform QM and CCA by significant margins. SPECD2 has the lowest MAE for spatial correlation in TMAX compared to the other methods; however, SPECD1 has slightly lower Wasserstein distance and in upper quantile MAE. Both SPECD methods have significantly lower cross correlations compared to QM and CCA.
For PRCP, SPECD still outperforms QM and CCA overall. However, SPECD2 has the highest MAE for the upper quantile than any of the competing methods, whereas SPECD1 has the lowest. We believe this is due to SPECD2 being the only method which explicitly models the spatial dependence structure. As a consequence, the marginal errors are somewhat higher. This is further exacerbated by PRCP being modeled on the log-scale, as errors get magnified when transformed back to the original scale. SPECD1, which only considers individual grid points, does not suffer from this issue. SPECD1 also has the lowest Wasserstein distance, but SPECD2 has the lowest MAE in spatial correlation estimates due to its underlying assumptions. Both SPECD algorithms perform comparably in their estimates of the proportion of zeros in PRCP data.
Overall, we see that SPECD comprehensively outperforms QM and CCE while calibrating the joint distribution of TMAX and PRCP over both the training and validation periods. The only exception lie with the heavy tails of PRCP, where SPECD runs into some issues when both the response and covariates are on the log scale. A table with the mean and standard errors for all the estimates, based on the 100 datasets, is provided in Appendix C. We also note minimal loss in predictive accuracy when going from SPECD1 to SPECD2 for marginal quantities. SPECD2 still outperforms QM and CCA, but not has the added benefit of incorporating spatial information in a theoretically justified manner.
@@ 4.1 Overview
We implemented SPECD for daily observational and GCM data from 1951–2014 on two separate $5\times 5$ grids as shown in Figure 1. We separated the data into a training period (1951–2000) and a validation period (2001–2014), and fitted the SPECD model (corresponding to SPECD2 from the numerical study) to the training period data. The Vecchia neighbor sets are selected using max-min ordering, which can result in significant improvements over other coordinate based orderings (Guinness, 2018).
@@
We model each month’s data separately to account for seasonality; for each month and at every location conditioned on its neighboring set of $m=10$ locations, we fitted SPQR models for TMAX and the transformed PRCP. Each SPQR model has several hyperparameters, which were chosen based on a grid search across different hyperparameter values (not presented). We choose each SPQR model to have 2 hidden layers with 30 and 20 nodes, and 20 output knots for both variables; the hidden layers have ReLU activation functions (Nair and Hinton, 2010). The batch size and learning rate is set to 100 and 0.001 respectively, and the SPQR model is run for 300 iterations. In each model, $20\%$ of the training data is used as a validation set, with early stopping enforced if the validation set loss fails to decrease for more than 5 iterations.
@@
The SPQR models (for each month, location, and variable) can be fit in parallel, since the Vecchia approximation decomposes the joint density of Y into independent univariate conditional densities. Predictions for every month can also be carried out independently, and therefore the projection/calibration step occurs in parallel. Within each month, however, predictions need to be made in the same sequence as the ordering of the Vecchia neighbor sets. Predictions for TMAX and PRCP were obtained at each location using Algorithm 2.
A key requirement for evaluating predictive models based on out-of-sample performance is for the training and validation data to arise from the same population distribution. In the context of density correction, the spatial dependence structure should therefore be identical in the training (1951–2000) and validation (2001–2014) periods. An exploratory study in Appendix D examines this behavior, by evaluating and plotting the spatial correlation for each pair of locations, during the training and validation periods. These are done separately for every month, in line with the density correction that is done separately for each month. Ideally, we want the differences in the spatial correlations between the two periods to be centered at zero, and have low variability. However, our study indicates that neither of these two properties hold for large sections of our data. Spatial correlations for PRCP show higher variability between the two periods than TMAX, possibly due to the presence of zeros in the data. The differences were often biased away from 0, indicating that the correlations have either strengthened/weakened over time. There were also situations where the model data and the observational data were biased in different directions; e.g., the spatial correlation in the model data decreases over time, while it strengthens in the observational data. Deviations were more frequent and more pronounced for pairs of locations with relatively low spatial correlations. While both regions demonstrated these patterns, it was especially prominent for the SW. We believe this to be at least in part due to the complex topography of the four-corners region. As a consequence, we anticipate that spatial correlations for the validation period might not be adequately calibrated based on the fitted models. SPECD is likely to be disproportionately affected compared to classical methods in such scenarios, since neural networks can struggle in generalizing outside the range of the observed data, or in the case of covariate drift. However, these issues are largely absent when we only consider in-sample predictions, i.e., when evaluating the density correction of data from 1951–2000. In situations with out-of-sample non-stationarity, SPECD can still realistically model spatial (and temporal) structures in the data, balancing flexibility and scalability. Section 5 discusses potential approaches to mitigate the effects of nonstationarity that we observe in the current study.
4.2 Evaluating density-corrected output based on SPECD
(a) Density curves for training period of 1951–2000.
(b) Density curves for validation period of 2001–2014.
@@ 4.2 Evaluating density-corrected output based on SPECD
Figure 5: Marginal density curves of climate model (Mod), observed (Obs), and calibrated (Cal) TMAX and PRCP data in the Southwest (SW) and Southeast (SE), pooled across location and month. PRCP results are presented on the log scale.
@@
In this section, we present results from density correction at both sets of locations for the training and validation periods. Figure 5 plots density curves of density corrected TMAX and log-scale PRCP for the SE and SW regions. Each curve pools the entire training/validation period data across all 25 locations, and we see that the calibrated TMAX and PRCP closely match the observed data for both time periods. Note that the density plots for PRCP are displayed in the transformed log scale since visual inspection of the density plots are most informative in the scale that the data was analyzed. However, all remaining plots and metrics for PRCP are presented in the raw data scale. Fitting separate monthly models for TMAX captures the pronounced bimodal distribution in the SW region with different modes for observed and model data. On the other hand, there are substantial differences between the observed and model densities of PRCP at both tails of their respective distributions. For both variables, SPQR is able to successfully calibrate the model data to match the density of the observational data.
(a) Monthly 0.95 quantiles for training period of 1951–2000.
(b) Monthly 0.95 quantiles for validation period of 2001–2014.
@@
Figure 6: Monthly 0.95 quantiles of TMAX and PRCP data in the Southwest (SW) and Southeast (SE). Each point corresponds to a particular month and location.
(a) Proportion of zeros in training period (1951–2000) for SW (left) and SE (right).
(b) Proportion of zeros in validation period (2001–2014) for SW (left) and SE (right).
@@
Figure 7: Monthly Proportion of zeros in PRCP data for the Southwest (SW) and Southeast (SE). Each point corresponds to a particular month and location.
@@
Figures 67 plot the 0.95 quantiles of TMAX and PRCP, and the proportion of zeros for PRCP, respectively. Each panel displays the effect of density correction by comparing it against the observational data for every location and month. For the upper 0.95 quantiles, the uncalibrated model values for TMAX have noticeable bias, something that can be visually verified in Figure 5. The calibrated quantiles virtually eliminates this bias while also having lower uncertainty than the uncalibrated data, for both training and validation periods. The density correction also improves upper quantile estimation for PRCP; however, there is noticeably more uncertainty in the PRCP estimates than TMAX. The calibration is adequate for the training period, but SPQR has difficulty calibrating the upper tail of the out-of-sample PRCP for the validation period. This is consistent with the numerical study in Section 3. We also note that the calibration is overall better for the SW than the SE. For estimating the proportion of zeros, we treat model precipitation below $0.001$ mm as zero precipitation, since it is not measurable by physical instruments like rain gauges. This was also done for the SPQR estimates since the continuous fitted distribution for PRCP will never produce exact zeros. The uncalibrated model data (red) in Figure 7 severely underestimates the proportion of zeros due to the drizzle effect. The calibrated density of PRCP is able to accurately represent the proportion of zeros that are observed in the real data. Calibration for the validation period (Fig 7(b)) has some bias indicative of an over-correction and more variability compared to the training period (Fig 7(a)), but the overall results suggest that our log-transform for modeling PRCP is adequate for this application.
(a) Cross correlations for training period of 1951–2000.
(b) Cross correlations for validation period of 2001–2014.
@@
Figure 8: Monthly cross-correlations between TMAX and PRCP in the Southwest (SW) and Southeast (SE). Each point corresponds to a particular month and location.
(a) Spatial correlations for training period of 1951–2000.
(b) Spatial correlations for validation period of 2001–2014.
Figure 9: Annual pairwise spatial correlations of TMAX and PRCP data in the Southwest (SW) and Southeast (SE). Each point corresponds to a pair of locations.
@@
Figure 8 plots the effect of density correction on the cross correlation between variables. The uncalibrated data has positive bias in the cross correlations for both regions, with the bias in the SW being more pronounced. Density correction was able to virtually eliminate the bias and provided highly accurate and precise estimates of the cross correlation during both the training and validation periods. This is in part because the conditional density of PRCP depends on both current and past TMAX values, and therefore SPQR models the dependence between the two variables.
Figure 9 shows pairwise spatial correlations for the calibrated and uncalibrated data. Each panel contains values for 300 pairs of locations. For the training period (Fig 9(a)), values are largely concentrated around the diagonal, suggesting the the calibrated data has similar spatial correlation compared to the observed data. The uncertainty in the estimates is higher for the SW than the SE, and pairs of locations with lower spatial correlations have more uncertainty in their estimates. For the validation period (Fig 9(b)), the uncertainties in the estimates are more pronounced. Even in such situations, the maximum error in the estimates is only $\approx 0.10$. The differences in the amount of calibration between the training and validation periods can be attributed to the dependence structure changing between the two periods and to modeling PRCP in the log-scale; additional details are provided in Appendix D.
(a) Monthly means for training period of 1951–2000.
(b) Monthly means for validation period of 2001–2014.
Figure 10: Monthly means of TMAX and PRCP data in the Southwest (SW) and Southeast (SE), pooled across all 25 locations.
Finally, we also verified whether SPECD can capture monthly and seasonal characteristics of PRCP and TMAX. Figure 10 plots the monthly means of the two variables for each region. For both the training (Fig 10(a)) and validation (Fig 10(b)), TMAX is calibrated very accurately, lining up with the observed data across all panels. The calibration process maps the distribution of the model data onto the distribution of the observational data; since the shape of the distribution of TMAX does not change between the two periods, the calibration works very well for the variable. The distribution of monthly PRCP, however, shows changes between the training and the validation periods. The changes occur in both the model and observational process, and are more pronounced in the SE compared to the SW. In the SW, the training period PRCP is calibrated adequately, with some deviations in the winter months of Dec-Feb. For the validation period, the calibration struggles in months where the relative behavior of model and observational data is very different from the training period. For example, Feb and Sep are two months where the model and observational data differ the most during the training period, but which are very close to each other in the validation period. A similar pattern is seen for SE PRCP. The calibration for the training period is adequate for most months, but the calibrated PRCP in the validation period is an interpolation of the training and validation period means of the observational data. As before, months where the relative behavior of the model and observed data change the most between the two periods are ones where the calibration struggles the most. Since neural networks often find generalizing outside the range of the training data challenging, changes in the joint distribution of observed and model data affect predictive accuracy of SPECD since it has not been trained on the distribution of PRCP in the validation period. Appendix E compares additional seasonal density fits, the upper quantiles of TMAX and PRCP, and proportion of zeros for PRCP. In particular, we note that though the calibration of monthly 0.95 quantiles of PRCP show the same behavior as the mean (likely due to the log-transform), the proportion of zeros are well-calibrated for both periods.
4.3 Comparison of SPECD with CCA and QM